OpenStudy (anonymous):

How do I write these in interval notation? {x element R : x<-1 or x>=1} {y element R : 0<=y<1 or y>1}

OpenStudy (mertsj):

\[(-\infty,-1)\cup[1,\infty)\]

OpenStudy (mertsj):

\[[0,1)\cup(1,\infty)\]

OpenStudy (anonymous):

Thank you!!

OpenStudy (mertsj):

yw